# Firstname1 Lastname1    Firstname2 Lastname2    Firstname3 Lastname3    Firstname4 Lastname4    Firstname5 Lastname5    Firstname6 Lastname6    Firstname7 Lastname7    Firstname8 Lastname8    Firstname8 Lastname8

### v1

@@ D.1 Fixed Modulus
Figure 18: Test accuracy vs token positions for $m=2048=2^{11}$, depth=1, various values of $a,c$. (a,b) $n_{heads}=1$ (c,d) $n_{heads}=4$. The accuracies for different $a$ and $c$ are exactly on top of each other.
Figure 19: Projections along Top 6 principle components of the embedding matrix, for $m=512$, depth=1, $n_{heads}=1,d_{model}=768$.
D.2 Unseen Modulus
Pruning heads corresponding to irrelevant prime factors
To further validate our findings from Figure 6 regarding the correlation between attention heads and digit-wise accuracy, we conduct additional pruning experiments. The results of these experiments are presented in Figures 20 and 21.
((a))
((b))
((c))
@@ D.2 Unseen Modulus
((a))
((b))
((c))
Figure 21: Per-digit accuracy for $m_{\mathrm{test}}=2352$ after pruning specific attention heads. (a) Results from the same model used in Section 4.3, identical to Figure 6 (b2); (b) Performance after pruning the attention head responsible for grouping numbers by their values modulo $3$, which is relavent for one specific digit in this case. After pruning, the model’s performance shows a clear degradation, with the strongest one happening exactly at the digit corresponding to modulo $3$ (c) Performance after pruning the attention head responsible for grouping numbers by their values modulo $2$. The model’s performance on $m_{\mathrm{test}}$ got obliterated, as there are many base-$2$ digits in the radix representation of $m_{\mathrm{test}}$ in this case.
Patching other heads
@@
((a))
((b))
((c))
((d))
@@
Evidence for Step iii
In Figure 23, we present attention patterns and token distance statistics for a selected attention head in later layers, with the model the same as the one used in Section 4.3. The token distances are measured by the spacing between keys with top-$4$ attention weights for a given query. Our analysis reveals that the model develops a head capable of dynamically adjusting their lookback distance for computations based on $m_{\mathrm{test}}$, which we interpret as evidence for step iii of the algorithm proposed in Section 4.3.
@@
Appendix E Training Variance
The model may converge to different solutions depending on the random seed used for initialization and the random seed for dataset generation. LABEL:{fig:med_fixp_appendix} shows the median value across five runs, with the shaded region representing the range between the minimum and maximum values. For larger moduli, not all models successfully find a solution that achieves $100\%$ test accuracy.
Figure 24: The median number of in-context sequence elements required for $100\%$ test accuracy across five runs. The min-max range is shown as a shaded region.
Appendix F Scaling Up the Modulus
F.1 Fixed Modulus
@@
For each modulus $m=2^{k}$, where $k\in[16,32]$, we train a 2-layer GPT model with an embedding dimension of 1024 and a vocabulary size 256. The train set consists of $n_{a}=1024$ multipliers and $n_{c}=1024$ increments, selected via the Hull-Dobell theorem. One LCG sequence of length 512 is included in the train set for each (a, c) pair, resulting in a total training set size of $n_{a}\times n_{c}=1,048,576$. For each modulus, the model is trained for 200,000 steps with a batch size of 512. The context length is 511×511$\times$511 × the digit length of $p$ in the byte representation. For example, when $m=2^{28}$, the digit length in byte representation is 4, while for $m=2^{23}$, the digit length is 3.
F.2 Unseen Modulus
@@
We train a 6-layer GPT model on a dataset that comprises $n_{m}=32,768$ moduli, with $n_{a}=128$ training multipliers and $n_{c}=1$ training increment per modulus. This results in a total of $n_{m}\times n_{a}\times n_{c}=4,194,304$ sequences of length 512 in the training set. The moduli are sampled from $m_{train}\in\{x\in\mathbb{Z}\mid 1024\leq x<65536\}$ in Fig. 25a, from $m_{train}\in\{x\in\mathbb{Z}\mid 1024\leq x<59049\}$. Multipliers are selected based on the Hull-Dobell theorem when sufficient qualifying $a$ values are available; otherwise, random $a$ values are used to ensure $128$ multipliers for each modulus. The models are trained on 4 million sequences for 2 million steps with a batch size of $128$. Because LCGs are typically defined for moduli that are powers of prime numbers, the model is tested on moduli that are powers of the primes 2, 3, 5, and 7. The test set consists of 512 unseen $a$ values and 64 unseen $c$ values for each test modulus.
@@
Test performance is influenced by the tokenization base, exhibiting a bias toward moduli that share the same base as the tokenization method. For instance, in Fig. 25 (a) when using a byte-level representation, the model achieves better performance on moduli $p=2^{k}$ compared to $p=3^{k}$, $5^{k}$, or $7^{k}$. As contrasted with Fig. 25 (b) where the tokenization base is $243=3^{5}$, the model performs better on moduli $p=3^{k}$. This behavior is likely due to the property of LCGs, where for moduli that are powers of a prime $b$, the lowest $k$-th digit exhibits a period of $b^{k}$. Tokenization in such a base highlights this periodic structure, making it more apparent and easier for the model to leverage during training and prediction.
((a))
((b))

### v2

@@ Appendix G Fixed Modulus Interpretability
Figure 25: Projections along Top 6 principal components of the embedding matrix, for $m=512$, depth=1, $n_{\mathrm{heads}}=1,d_{\mathrm{model}}=768$.
Figure 26: Projections along Top 6 principal components of the embedding matrix, for $m=7776$, depth=1, $n_{\mathrm{heads}}=1,d_{\mathrm{model}}=768$.
Figure 27: Projections along Top 6 principal components of the embedding matrix, for $m=1800$, depth=1, $n_{\mathrm{heads}}=1,d_{\mathrm{model}}=768$.
Appendix H Unseen Modulus Interpretability
H.1 Extra PCA plots for first layer heads
We present additional PCA analyses of attention heads shown in Figure 9, examining $\mathrm{PCA}(\bm{H}^{(h)}[:,t,:])$ for each head $h$ across various combinations of $a$, $c$, $x_{0}$, $m_{\mathrm{test}}$, and position $t$. The results are depicted in Figure 28, where we include two additional first-layer heads (heads 2 and 3) that are responsible for performing operations related to modulo $5$. Notably, with different choices of $a$, $c$, $m_{\mathrm{test}}$ and $t$ compared to Figure 9 (a1, b1, c1), the clustering behavior remains unchanged.
Note the emergence of attention heads dedicated to processing modulo $2$, $3$, $5$, and $7$ can likely be attributed to the prevalence of these prime factors in the training set.
(a) $a=13$, $c=3$, $m_{\mathrm{test}}=2048$, $t=0$
(b) $a=9$, $c=3$, $m_{\mathrm{test}}=2048$, $t=128$
(c) $a=61$, $c=7$, $m_{\mathrm{test}}=1800$, $t=47$
Figure 28: PCA analysis of first-layer heads. Although the ordering of numbers varies, the grouping behavior discussed in Section 4.3 remains invariant to changes in $a$, $c$, $m$, $x_{0}$, and $t$. Note that the last head is the head that appears in Figure 10, which does not exhibit a strong grouping bias.
H.2 Pruning heads corresponding to irrelevant prime factors
To further validate our findings from Figure 9 in Section 4.3 regarding the correlation between attention heads and digit-wise accuracy, we conduct additional pruning experiments. The results of these experiments are presented in Figures 29 and 30.
(a) Original Model
(b) Prune modulo $3$ head
(c) Prune modulo $14$ head
@@ H.2 Pruning heads corresponding to irrelevant prime factors
(a) Original Model
(b) Prune modulo $3$ head
(c) Prune modulo $2$ head
Figure 30: Per-digit accuracy for $m_{\mathrm{test}}=2352$ after pruning specific attention heads. (a) Results from the same model used in Section 4.3, identical to Figure 9 (b2); (b) Performance after pruning the attention head responsible for grouping numbers by their values modulo $3$, which is relevant for one specific digit in this case. After pruning, the model’s performance shows a clear degradation, with the strongest one happening exactly at the digit corresponding to modulo $3$ (c) Performance after pruning the attention head responsible for grouping numbers by their values modulo $2$. The model’s performance on $m_{\mathrm{test}}$ got obliterated, as there are many base-$2$ digits in the RNS representation of $m_{\mathrm{test}}$ in this case.
H.3 Patching other heads
@@
(a) layer 1, head 4
(b) layer 2, head 4
(c) layer 3, head 6
(d) layer 4, head 1
@@
H.4 In-accurate estimation of $m_{\mathrm{test}}$
From the cosine-similarity panel in Figure 10, we observe that the model’s estimation approximates the target $m_{\mathrm{test}}=2048$. However, detailed analysis reveals that the highest cosine-similarity occurs at $m_{\mathrm{est}}=2033=19\cdot 107$, with neighboring values exhibiting similarly high cosine-similarity values. If we assume $m_{\mathrm{est}}$ represents the model’s internal belief, then the prime representation would consist solely of powers of $19$ and $107$. Such a representation can only produce periodic structures for the $k$-th bit of a binary number when $r=\operatorname{lcm}(19,2^{k})$. Consequently, patterns before reaching that period would appear random in this representation, providing a weaker signal compared to the correct representation. This explains why the model preferentially selects binary representation for lower bits when $m_{\mathrm{test}}=2048$.
For higher bits, the low-bit representation can be determined up to $2^{k}$ bits through copying. If the model utilizes this information in later stages, the precision of $m_{\mathrm{est}}$ can be drastically improved. Specifically, when $\lfloor{m_{\mathrm{est}}}/2^{k}\rceil=\lfloor{m_{\mathrm{test}}}/2^{k}\rceil$, the model can identify the correct representation as long as $|m_{\mathrm{test}}-m_{\mathrm{est}}|<2^{k}$. This argument can be extended to any other composite $m_{\mathrm{test}}$. Note that this argument is hypothetical; further proof of this mechanism remains in future work.
H.5 Evidence for Step iii
In Figure 32, we present attention patterns and token distance statistics for a selected attention head in later layers, with the same model as the one used in Section 4.3. The token distances are measured by the spacing between keys with top-$4$ attention weights for a given query. Our analysis reveals that the model develops a head capable of dynamically adjusting their lookback distance for computations based on $m_{\mathrm{test}}$, which we interpret as evidence for step iii of the algorithm proposed in Section 4.3.
@@ H.5 Evidence for Step iii
Appendix I Scaling Up the Modulus
I.1 Base-b tokenization
To convert an integer $x$ to base $b$ with the least significant digit (LSD) first, repeatedly divide $x$ by $b$, storing remainders:
$d_{k}=x\mod b,\quad x=\lfloor x/b\rfloor.$ (17)
Stop when $x=0$. The sequence $(d_{0},d_{1},\dots)$ is the base-$b$ representation in LSD-first order.
Example: Converting $x=3,214,748,365$ to base 256:
$3,214,748,365\div 256$ $=12,557,610\text{ remainder }205,\quad d_{0}=205,$
$12,557,610\div 256$ $=49,053\text{ remainder }42,\quad d_{1}=42,$
$49,053\div 256$ $=191\text{ remainder }157,\quad d_{2}=157,$
$191\div 256$ $=0\text{ remainder }191,\quad d_{3}=191.$
The final LSD-first representation consists of four tokens:
$(205,42,157,191)_{256}.$
Each token is then one-hot encoded into a $b$-dimensional vector, where only the index corresponding to the token value is set to 1. These one-hot vectors are then fed into a token embedding layer with an embedding dimension of $d_{model}=1024$.
I.2 Abacas Embeddings
The positional embedding for the $j$-th lower digit of the $i$-th number in the sequence is defined as:
$\mathrm{PosEmbed}(T_{i,j})=E_{\mathrm{int}}(i)+E_{\mathrm{digit}}(j)$ (18)
where $T_{i,j}$ represents the token corresponding to the $j$-th digit of the $i$-th integer.
  • $E_{\mathrm{int}}(i),E_{\mathrm{digit}}(j)\in\mathbb{R}^{d_{\text{model}}}$ are learnable embeddings.
  • $E_{\mathrm{int}}(i)$ encodes the integer’s position in the sequence.
  • $E_{\mathrm{digit}}(j)$ encodes the digit’s relative position within the integer.
This embedding scheme ensures that each token captures both the integer’s global position and the byte’s local position. Figure 11 provides a visualization of base-$b$ tokenization and abacus embedding.
I.3 Fixed Modulus
@@
For each modulus $m=2^{k}$, where $k\in[16,32]$, we train a 2-layer GPT model with an embedding dimension of 1024 and a vocabulary size 256. The train set consists of $n_{a}=1024$ multipliers and $n_{c}=1024$ increments, selected via the Hull-Dobell theorem. One LCG sequence of length 512 is included in the train set for each (a, c) pair, resulting in a total training set size of $n_{a}\times n_{c}=1,048,576$. For each modulus, the model is trained for 200,000 steps with a batch size of 512. The context length is 512×512$\times$512 × the digit length of $m$ in the byte representation $-1$. For $m=2^{32}$, the digit length in byte representation is 4; therefore, the context length is 2047. Training was performed using 4 A100 GPUs over a total duration of 21.82 hours. For $m=65536$, the digit length in byte representation is 2, resulting in a context length of 1023, with training taking 4.83 hours. For each modulus, the test set includes 512 unseen $a$ values and 64 unseen $c$ values selected via the Hull-Dobell theorem.
The model may converge to different solutions depending on two random seeds: one for model initialization and batch shuffling, and another for dataset generation. LABEL:{fig:med_fixp_appendix} shows the median performance across five runs, with the shaded region representing the range between the minimum and maximum values. For larger moduli, not all models successfully find a solution that achieves $100\%$ test accuracy.
(a) PyTorch seed = 9, dataset seed = 71
(b) PyTorch seed = 10, data seed = 71
(c) PyTorch seed = 11, data seed = 71
(d) PyTorch seed = 11, data seed = 71
(e) PyTorch seed = 11, data seed = 72
(f) PyTorch seed = 11, data seed = 73
Figure 33: Test accuracy vs. Number index for $m=2^{20}$. First Row: Three models trained on the same dataset, each using a unique PyTorch random seed that controls model initialization and batch shuffling. Second Row: Three models trained on different datasets, with each dataset generated using a unique random seed controlling NumPy randomness for sampling $a$, $c$, and $x_{0}$. All models converged to solutions that achieved and sustained 100% test accuracy, but differed in the number of in-context examples required to reach this performance.
(a) PyTorch seed = 9, dataset seed = 71
(b) PyTorch seed = 10, dataset seed = 71
(c) PyTorch seed = 11, dataset seed = 71
(d) PyTorch seed = 11, dataset seed = 71
(e) PyTorch seed = 11, dataset seed = 72
(f) PyTorch seed = 11, dataset seed = 73
Figure 34: Test accuracy vs Number index for $m=2^{32}$. First Row: Three models trained on the same dataset, each using a unique PyTorch random seed that controls model initialization and batch shuffling. Second Row: Three models trained on different datasets, with each generated using a unique random seed controlling NumPy randomness for sampling $a$, $c$, and $x_{0}$. Only one of the five models found a solution that achieved and sustained 100% test accuracy.
Figure 35: Median number of in-context sequence elements required to achieve 100% test accuracy across five runs. The shaded region represents the min-max range.
I.4 Unseen Modulus
@@
We train a 6-layer GPT model on a dataset that comprises $n_{m}=32,768$ moduli, with $n_{a}=128$ training $a$ values and $n_{c}=1$ training $c$ values per modulus. This results in a total of $n_{m}\times n_{a}\times n_{c}=4,194,304$ sequences, each of length 512 in the training set. In Figure 36a, where the tokenization base is 256, $655361024<m_{\mathrm{train}}<$ 65536. In Figure 36b, where the tokenization base is 243, $590491024<m_{\mathrm{train}}<$ 59049. Multipliers are selected based on the Hull-Dobell theorem when sufficient qualifying $a$ values are available; otherwise, random $a$ values are used to ensure $128$ multipliers for each modulus. The models with approximately 76M parameters were trained on 16 million sequences over 400,000 steps, using a batch size of 128 on a single H100 GPU for 22.62 hours. Because LCGs are typically defined for moduli that are powers of prime numbers, the model is tested on moduli that are powers of the primes 2, 3, 5, and 7. The test set consists of 512 unseen $a$ values and 64 unseen $c$ values selected via the Hull-Dobell theorem for each test modulus.
@@
Test performance is influenced by the tokenization base, exhibiting a bias toward moduli that share the same base as the tokenization method. For instance, in Figure 36 (a) when using a byte-level representation, the model achieves better performance on moduli $m_{\mathrm{test}}=2^{k}$ compared to $m_{\mathrm{test}}=3^{k}$, $5^{k}$, or $7^{k}$. As contrasted with Figure 36 (b) where the tokenization base is $243=3^{5}$, the model performs better on moduli $m=3^{k}$. This behavior is likely due to the property of LCGs, where for moduli that are powers of a prime $b$, the lowest $k$-th digit exhibits a period of $b^{k}$. Tokenization in such a base highlights this periodic structure, making it more apparent and easier for the model to leverage during training and prediction.
(a) Tokenization base = $256=2^{8}$
(b) Tokenization base = $243=3^{5}$